Calculating the Predetermined Overhead Rate, Applying Overhead to Production, Reconciling Overhead at the End of the Year, Adjusting Cost of Goods Sold for Under- and Overapplied Overhead

At the beginning of the year, Han Company estimated the following:

Overhead $210,000
Direct labor hours 70,000

Han uses normal costing and applies overhead on the basis of direct labor hours. For the month of January, direct labor hours were 8,150. By the end of the year, Han showed the following actual amounts:

Overhead $216,000
Direct labor hours 69,600

Assume that unadjusted Cost of Goods Sold for Han was $276,000.

Required:

1. Calculate the predetermined overhead rate for Han. Round your answers to the nearest cent, if rounding is required.
$ per direct labor hour

2. Calculate the overhead applied to production in January. (Note: Round to the nearest dollar, if rounding is required.)
$

3. Calculate the total applied overhead for the year.
$

Was overhead over- or underapplied? By how much?
Underapplied  overhead $

4. Calculate adjusted Cost of Goods Sold after adjusting for the overhead variance.
$
Answer #1

1) Predetemine overhead rate = Estimated overhead/estimated labor hour = 210000/70000 = $3 per labor hour

2) Overhead applied in january = 8150*3 = $24450

3) Total overhead applied for the year = 69600*3 = $208800

Actual overhead = $216000

Under applied overhead = 216000-208800 = $7200

4) Adjusted cost of goods sold = 276000+7200 = $283200